# AtomSurf: Surface Representation for Learning on Protein Structures

### v3

@@ 4.2 Synergy in Combined Representations
Our method outperforms single modalities in all three tasks, while adhering to the Atom3d benchmark protocol. This result highlights the synergy between the two representations. Achieving this is noteworthy because, fist, the input features remain minimal, and second, to maintain a consistent parameter count, both the graph and surface encoders are considerably condensed. An interesting result is that even in the case of PSR, where surfaces do not intuitively seem relevant, the mixed model outperforms its graph counterpart with a comfortable margin. One possible interpretation for this result is DiffusionNet’s ability to perform long-range message passing.
In addition, we compare to popular, state-of-the-art models outside of the benchmark constraints. Namely, we compare against ProNet (Wang et al., 2022a) and GearNet (Zhang et al., 2023), which are graph-based methods at the residue level, but encode the local geometry of residues as features; and to the atomic-level extension of GVP (Jing et al., 2021). Finally, we include their extension with features derived from protein language models (Zhang et al., 2023). Please note that differing parameters and training procedures do not necessarily follow the protocol of the original benchmark. For this reason, we also include the performance of our second proposed model, AtomSurf, which uses all input features and more recent encoders. The results are presented in Table 2.
@@
Pretrained/ PIP MSP PSR
Parameters Auroc Auroc Local R Global R
ProNet (Wang et al., 2022a) No/2M 0.871 - 0.632 0.849
GVP (Jing et al., 2021) Yes/11M 0.866 0.68 0.511 0.845
GVP-ESM (Zhang et al., 2023) Yes/11M - 0.617 - 0.866
GearNet-ESM (Zhang et al., 2023) Yes/42M - 0.685 - 0.829
AtomSurf-bench No/200k 0.876 0.707 0.452 0.831
AtomSurf No/600k 0.907 0.715 0.613 0.861
Our first result is that even AtomSurf-bench is competitive with more highly engineered approaches and outperforms them in two of the three evaluated tasks. The PSR task, which is aimed at detecting non-canonical protein structures, appears to benefit methods that integrate explicit biological insights into protein geometry, such as the computation of side-chain angles.
Moreover, AtomSurf successfully results in another performance boost on this benchmark, setting a new clear state-of-the-art performance on the PIP and MSP tasks, and closing the gap on PSR. Importantly, our method does not include any pretraining and uses less than half the number of parameters than any other method.
4.3 Evaluation of binding site detection
In addition to the Atom3d benchmark tasks, we evaluate our approach on the task of protein binding sites prediction. Binding sites are regions of protein surfaces where a partner (that can be another protein, an RNA molecule, a small molecule) interacts with the protein. A fine characterization of those regions helps in understanding protein functions and plays a significant role in drug design and discovery. In this section we focus on binding site prediction in the context of immunology: predicting the interaction between an antibody and its target, denoted as the antigen. Improved understanding and ability to predict such interactions has direct applications in antibody-based treatments, which represent a highly promising therapeutic avenue (Kaplon et al., 2023; Jamali et al., 2024).
Pegoraro et al. (2024) introduced a dataset containing 235 antibody–antigen complexes, with 186 for training and 49 for testing. The task consists in predicting the the residues involved in the binding site on the antibody and antigen separately. Importantly, the network has access to both structures to perform its prediction, but not to how they interact. The performance on the antibody side is computed only on the Complementarity Determining Regions (CDRs). The method proposed in their article, MIXmean-egnn PiNet, is a relevant comparison to ours, because it averages the prediction made by a surface model, and the ones obtained with a graph-based model. We denote this method as GEP, and present the results in Figure 4 and Table 4.
@@ 4.3 Evaluation of binding site detection
Figure 3: Performance of our model on the binding site detection task.
Figure 4: Performance of our hybrid approach on the MaSIF-ligand task.
As can be seen in our results, AtomSurf-bench has a similar performance on the antigen binding site prediction, but outperforms other methods with a comfortable margin on the antibody side (+0.19 MCC). An accurate prediction of the CDR residues involved in the recognition of an antigen is key for antibody specificity optimization.
4.4 Evaluation of Binding Sites Classification
We also evaluate our proposed approach on the task of ligand-binding preference prediction for protein binding sites, introduced in (Gainza et al., 2020). It holds 1459 protein structures bound to one of the seven most common ligands in the PDB. Given a binding site, the task amounts to predicting its corresponding co-factor. We benchmark our approach against MaSIF (Gainza et al., 2020) and HMR (Wang et al., 2023), with the latter being recognized as state-of-the-art on this task. Additionally, considering the notable performance of ProNet on the previous benchmark, we include it in this experiment. We use balanced accuracy (the average recall achieved for each class) as our performance metric consistent with prior studies.
4.5 Qualitative Results
To visualize the predictions of our binding site model, we display the interaction probability predicted by our model across two protein surfaces, a homodimer and a heterodimer, an plot the results in Figure 5. Despite minor prediction errors, such as the misidentification of residues in the lower region of 3dbh, our results clearly show that the model effectively identifies binding sites on proteins.
@@ 4.5 Qualitative Results
In the supplementary (Figures 6 and 7) we provide visualizations showing both the relation between the effective receptive field and diffusion times, as well as the distribution of learned times on the MSP task. These visualizations demonstrate that long-range communication enabled by learned diffusion is indeed useful, and can contribute to overall performance, thus shedding light on the relative strengths of surface-based learning.
@@ 4.6 Ablation Study
Finally, we examine the impact of different design choices on our tasks. All of these results are obtained in benchmark settings, both in terms of features (only atom-type as input) and of graph encoders (GCNs). In Appendix C, we introduce the sequential and bipartite scenario, which amount to variations in the connectivity of the blocks. In particular, the HoloProt method proposed in Somnath et al. (2021) falls into the sequential framework with just one encoding block and a MeshCNN (Hanocka et al., 2019) for protein encoder. For fairness, we replace MeshCNN with our encoder and refer to this setup as HoloProt.
Another major design choice is the choice of the Message Passing (MP) component. We explore the use of three possible message-passing networks. Motivated by the success of DGCNN (Wang et al., 2019), in our Att. setting, we discard the geometric notion of a neighborhood, allowing for potentially long-distance message passing. In this setting, all nodes from the graphs attend to all vertices from the surface. To deal with the incurred computational burden, we use the recent memory-efficient Flash Attention (Dao et al., 2022). We also explore the use of more conventional Graph Convolutional Networks (GCN setting) (Kipf & Welling, 2016) and the use of Graph ATtention networks (Veličković et al., 2017; Brody et al., 2021) for our final GAT setting. Our results are presented in Table 3.
Table 3: Ablation study of our method: We compare various architectural designs, and message-passing methods for our task.
MSP PIP PSR
Method MP AuROC AuROC local R global R
sequential - 0.609 0.855 0.319 0.71
HoloProt - 0.537 0.824 0.383 0.715
Att. 0.689 0.791 0.402 0.792
bipartite GCN 0.697 0.868 0.421 0.797
GAT 0.707 0.876 0.452 0.833
@@ 5 Conclusion & Limitations
Our work strongly supports the notion that leveraging multiple representations with unique strengths is a promising strategy for advancing protein analysis. Moreover integrating biological priors both within the learning frameworks and into information sharing across representations seems crucial for achieving high accuracy in challenging scenarios. Despite our optimizations, one of the current limitations of our approach is its significant memory requirements, highlighting the need for computationally more efficient surface-based pipelines, in line with suggestions by Sverrisson et al. (2021). Beyond addressing these challenges, our work can help pave the way to more powerful integrated multi-modal solutions for additional tasks within structural bioinformatics, including generative modeling, but also investigating ways in which information sharing across both representations and tasks can lead to improvements in robustness and accuracy.
@@
More specifically, we first provide additional data illustrations in Appendix A. Then, we provide a proof for Proposition 3.1 in Appendix B. Following this, detail the specifics of our implementation in Appendix C. Then, we provide an analysis of the computational aspects of our technique in Appendix D. Finally, we provide additional results in Appendix F.
@@ Appendix A Additional data analysis
In Figure 6, we plot the results of diffusing a Dirac initial distribution for several diffusion times, over the surface of the MDM2 protein, involved in apoptosis and cancer treatment. On the figure we see its groove, which corresponds to its binding site. Orange vertices correspond to the ones that have the highest probabilities, such that their sum represents 90% of the probability mass. A diffusion time of ten seems to convey a relevant scale for this binding site.
Figure 6: Visualization of a Dirac delta function $\delta_{x}$ at a point, diffused for several diffusion times.
In Figure 7, we plot the diffusion times learnt by a network (on the MSP task). We see that the network keeps a variety of scales, and has learnt to use large scales, effectively enabling long-distance message passing.
Figure 7: Histogram of the diffusion times obtained after training.
Appendix B Proof of Proposition 3.1
@@
Suppose $M_{2}$ is another mesh that is a scaled version of $M_{1}$ by a scaling factor $a$. Then, according to Bronstein & Kokkinos (2010):
@@
The heat kernel can be computed as per Sun et al. (2009):
@@ Appendix C Implementation Details
However, the sequential approach does not simultaneously leverage the distinct scales (local and global) offered by graph and surface processing. To overcome this, the bipartite approach processes graph and surface features concurrently in separate encoders, and merge them with a message passing incorporating learnable parameters, based on the equation: $\mathcal{X}_{out}=\alpha\mathcal{X}+\text{MP}_{gs}(\mathcal{H})+\text{MP}_{sg}(\mathcal{H})$. Those architectures are illustrated in Figure 8.
Figure 8: Illustration of our approach integrating surface and graph information.
@@ D.1 Complexity analysis
The spectral operation component ($k^{(1)}_{\text{diff}}$) does not depend strongly, but it has a strong dependency on the size of the mesh, which is not well addressed by reducing the number of parameters. Empirically, for the latent dimensions considered, this spectral component was the time bottleneck and thus, coarsening the graphs resulted in significant speedups. We adjusted the number of vertices to make the first term of similar magnitude to the others (obtained for a 0.1 coarsening rate). In this regimen, when dividing the parameter count equally between the three learnt components, we observed that the ProNet execution was the limiting operation, and hence conclude that our method has a similar throughput to graph-based method for a fixed parameter count. Moreover, we observed that our method seemed to require less parameters overall, and note that the parameter count is split between components, effectively resulting in smaller ProNets than in the pure ProNet network of the benchmark, and hence a slightly faster runtime overall.
@@ D.2 Memory and batching
In addition to individual objects being large, the large variance in the object size (see Figure 10) also prevents us from using large batch sizes. Indeed, due to the random composition of the batch, some unlucky batches contain several particularly large proteins, causing spikes in memory usage and out-of-memory errors. We validate in Figure 10 that using coarsened surfaces, we stay in the linear dependency on the number of vertices for the hidden sizes at stake.
@@
Figure 9: Distribution of the number of vertices in a dataset.
Figure 10: Memory consumption plotted as a function of the number of vertex, in a log-log scale.
This is illustrated in Figure 11, for which we tracked the memory consumption of our model through one epoch of learning. Observe the memory peaks that can result in over 20% excess memory consumption, keeping in mind that this is only one run over one epoch. When training several models for hours, the probability that one outlier results in a 50% increase forces user caution.
Figure 11: Memory consumption monitoring through an epoch, for different batch sizes.

### v4

@@ 4.2 Synergy in Combined Representations
Our method outperforms single modalities in all three tasks, while adhering to the Atom3d benchmark protocol. This result highlights the synergy between the two representations. Achieving this is noteworthy because, first, the input features remain minimal, and second, to maintain a consistent parameter count, both the graph and surface encoders are considerably condensed. An interesting result is that even in the case of PSR, where surfaces do not intuitively seem relevant, the mixed model outperforms its graph counterpart with a comfortable margin. One possible interpretation for this result is DiffusionNet’s ability to perform long-range message passing.
In addition, we compare to popular, state-of-the-art models outside of the benchmark constraints. Namely, we compare against ProNet (Wang et al., 2022a) and GearNet (Zhang et al., 2023), which are graph-based methods at the residue level, but encode the local geometry of residues as features; and to the atomic-level extension of GVP (Jing et al., 2021). Finally, we include their extension with features derived from protein language models (Zhang et al., 2023). Please note that differing parameters and training procedures do not necessarily follow the protocol of the original benchmark. For this reason, we also include the performance of our second proposed model, AtomSurf, which uses all input features and more recent encoders. Out of fairness, we follow exactly the training protocol and problem formulation for each considered task, and use externally reported performance for all other methods. The results are presented in Table˜4.
@@
PLM Params PIP MSP PSR
Auroc Auroc $R_{l}$ $R_{g}$
ProNet 2M 87.1 - 63.2 84.9
GVP 11M 86.6 68.0 51.1 84.5
+ pretraining 11M 87.4 71.1 51.5 84.8
GVP-ESM 11M - 61.7 - 86.6
GearNet-ESM 42M - 68.5 - 82.9
+ pretraining 42M - 70.2 - 86.3
AtomSurf-bench 200k 87.6 70.7 45.2 83.1
AtomSurf 600k 90.9 71.6 61.7 85.7
Our first result is that even AtomSurf-bench is competitive with more highly engineered approaches and outperforms them in two of the three evaluated tasks. It even outperforms pretrained methods on the PIP task. The PSR task, which is aimed at detecting non-canonical protein structures, appears to benefit methods that integrate explicit biological insights into protein geometry, such as the computation of side-chain angles.
Moreover, AtomSurf successfully results in another performance boost on this benchmark, setting a new clear state-of-the-art performance on the PIP and MSP tasks, and closing the gap on PSR. Importantly, our method does not rely on pretraining and uses less than half the number of parameters than any other method.
4.3 Evaluation on protein interaction predictions
In addition to the Atom3d benchmark tasks, we evaluate our approach on the task of protein binding sites prediction. Binding sites are regions of protein surfaces where a partner (that can be another protein, an RNA molecule, a small molecule) interacts with the protein. A fine characterization of those regions helps in understanding protein functions and plays a significant role in drug design.
Masif-ligand
We start by evaluating our proposed approach on the task of ligand-binding preference prediction for protein binding sites, introduced in (Gainza et al., 2020). It holds 1459 protein structures bound to one of the seven most common ligands in the PDB. Given a binding site, the task amounts to predicting its corresponding co-factor. We benchmark our approach against MaSIF (Gainza et al., 2020) and HMR (Wang et al., 2023), with the latter being recognized as state-of-the-art on this task. Additionally, considering the notable performance of ProNet on the previous benchmark, we include it in this experiment. We use balanced accuracy (the average recall achieved for each class) as our performance metric consistent with prior studies.
Antibody epitope prediction
Improved understanding and ability to predict the interaction between an antibody and its target, denoted as the antigen, has direct applications in antibody-based treatments, which represent a highly promising therapeutic avenue (Kaplon et al., 2023; Jamali et al., 2024). Pegoraro et al. (2024) introduced a dataset containing 235 antibody–antigen complexes. The task consists in predicting interacting residue on the antibody and antigen, from their two structures. The method proposed in their article, denoted as GEP, is a relevant comparison to ours, because it averages predictions made by a surface model and by a graph-based model. We present the results in Figure˜5.
@@ 4.3 Evaluation on protein interaction predictions
[…]
As can be seen in our results, AtomSurf has a comparable performance to GEP on the antigen binding site prediction, but outperforms it with a comfortable margin on the antibody side (+0.25 MCC points). An accurate prediction of the antibody residues involved in the antigen recognition is key for antibody specificity optimization.
Validation on PINDER
In addition to this relatively small dataset, we validate our approach on the recently proposed, large-scale PINDER dataset (Kovtun et al., 2024). We used the clustered version of this dataset that holds around 42k structures. Systems are split rigorously, based on interface similarity. Moreover, test systems are also available in their unbound form (Apo setting) and as predicted by AlphaFold2 (AF2 setting), representing a more realistic use case. We consider a first task formulated as PIP (Pinder-Pair) of predicting interacting pairs of residues and another close to Masif-Site (Gainza et al., 2020), only taking one protein as input to predict interacting residues (Pinder-site). We compare to ProNet in Table˜8, and include a comparison of accuracies in Supplementary Table˜4.
Table 8: Auroc of our method compared to ProNet on the PINDER dataset.
Task Pinder-Pair Pinder-Site
Split Holo Apo AF2 Holo Apo AF2
Pronet 80.1 78.2 73.5 74.3 70.7 60.6
AtomSurf 92.8 88.4 87.1 88.3 84.2 82
AtomSurf widely outperform the ProNet baseline on all tasks, splits and metrics. The performance over the apo set is decreased compared to the holo set, as well as the performance on predicted structures, as expected (Huang et al., 2024). However, our network is able to retain a remarkable accuracy across the board, validating its robustness.
4.4 Additional Results
Visualization of our predictions
We display the interaction probability predicted by our model across two protein surfaces, a homodimer and a heterodimer, and plot the results in Figure˜6. Despite minor prediction errors, such as the misidentification of residues in the lower region of 3dbh, our results clearly show that the model effectively identifies binding sites on proteins.
@@ 4.4 Additional Results
In the Supplementary Fig. 2 and 3, we provide visualizations showing both the relation between the effective receptive field and diffusion times, as well as the distribution of learned times on the MSP task. These visualizations demonstrate that long-range communication enabled by learned diffusion is indeed useful, and can contribute to overall performance, thus shedding light on the relative strengths of surface-based learning.
Assessment of the reproducibility
Supplementary Table˜3 shows the standard error of our estimate of the mean performance of AtomSurf, assessing its reproducibility and the significance of its difference. Across tasks, the performance gap is higher than the standard error.
@@ 4.5 Ablation Study
Model ablations
Finally, we examine the impact of different design choices on our tasks. In Appendix˜C, we introduce the sequential and bipartite scenario, which amount to variations in the connectivity of the blocks. Another major design choice is the choice of the Message Passing (MP) component. We explore several options, including discarding the geometric notion of a neighborhood, allowing for potentially long-distance message passing (Att. setting), as well as varying the number of blocks. Our detailed results are presented in Section˜F.4.
The sequential strategy displays underwhelming results (70.7 vs 60.9 on MSP for example), which could root from its incapacity to handle multi-scale, simultaneous message passing. Among the bipartite settings, Att. is consistently outperformed by the localized message-passing networks. In short, the best setting leverages several blocks of message passing, with enhanced results for the most interconnected networks. In addition, the geometry of the bipartite graph is important, and an attentive mechanism generally benefits the optimal mixing.
ESM embeddings ablation
Due to the strong performance of protein language models, we also perform an ablation by removing the ESM embeddings from our graph initial node features. Ablated model’s performance decreases as expected, showing that ESM is truly an important feature. However, even without ESM, our method retains state-of-the-art performance on most tasks, notably on interaction tasks such as Masif-Ligand (84 vs 81), PINDER (overall +10 AuROC points) and others (detailed results in Section˜F.5). Interestingly, on the PINDER dataset that has stricter splits, the performance drop seems to be reduced. One possible explanation for this result would be that those sequence-derived features contribute to memorization, whereas learned structural properties lead to better generalization. We believe that a more in-depth investigation of this phenomenon is an interesting direction for future work.
@@ 5 Conclusion & Limitations
Our work strongly supports the notion that leveraging multiple representations with unique strengths is a promising strategy for advancing protein analysis. Moreover, integrating biological priors both within the learning frameworks and into information sharing across representations seems crucial for achieving high accuracy in challenging scenarios. A validation on real-life scenarios is our next step to fully establish this method. Despite our optimizations, one of the current limitations of our approach is its significant memory requirements, highlighting the need for computationally more efficient surface-based pipelines, in line with suggestions by Sverrisson et al. (2021). Beyond addressing these challenges, our work can help pave the way to more powerful integrated multi-modal solutions for additional tasks within structural bioinformatics, including generative modeling, but also investigating ways in which information sharing across both representations and tasks can lead to improvements in robustness and accuracy.
Acknowledgements & Funding
The authors would like to thank Emanuele Rodolà and Marco Pegoraro for discussions that contributed to initiate this project. They also would like to thanks Erkan Turan for trying to setup more tasks, as well as Victor Laigle for making the Supplementary Figures on batch size variability.
V.M., S.A. and M.O. are supported by the ANR Chair AIGRETTE, the ERC Starting Grant No. 758800 (EXPROTEA) and the ERC Consolidator Grant No. 101087347 (VEGA). V.M. was additionally supported by DataIA and Sanofi. Y.M and B.C. are supported by Swiss National Science Foundation grants 310030_197724, TMGC-3_213750 and 200020_214843. This work was performed using HPC resources from GENCI–IDRIS (Grant 2023-AD010613356) and CITAS at EPFL.
Broader impact
This paper presents work to advance the field of machine learning on protein structure. There is no direct ethical or societal implication of this work, but potential indirect ones, none of which we feel must be specifically highlighted here.
@@
More specifically, we first provide additional data illustrations in Appendix˜A. Then, we provide a proof for Proposition˜3.1 in Appendix˜B. Following this, we detail the specifics of our implementation in Appendix˜C. Then, we provide an analysis of the computational aspects of our technique in Appendix˜D. Finally, we provide additional results in Appendix˜F.
@@ Appendix A Additional data analysis
In Supplementary Figure˜1, we illustrate the different representations that exist for protein structures.
Supplementary Figure 1: Diverse mathematical objects used to represent a protein structure, sequences, molecular surfaces (blue), atom-level and residue-level point clouds (red) and graphs (green). Effective machine learning for protein structures hinges on selecting the appropriate mathematical representation along by a compatible machine-learning technique.
A.1 Assessment of the scale to use
In Figure˜2, we plot the results of diffusing a Dirac initial distribution for several diffusion times, over the surface of the MDM2 protein, involved in apoptosis and cancer treatment. On the figure we see its groove, which corresponds to its binding site. Orange vertices correspond to the ones that have the highest probabilities, such that their sum represents 90% of the probability mass. A diffusion time of ten seems to convey a relevant scale for this binding site.
Supplementary Figure 2: Visualization of a Dirac delta function $\delta_{x}$ at a point, diffused for several diffusion times.
In Figure˜3, we plot the diffusion times learnt by a network (on the MSP task). We see that the network keeps a variety of scales, and has learnt to use large scales, effectively enabling long-distance message passing.
Supplementary Figure 3: Histogram of the diffusion times obtained after training.
Appendix B Proof of Proposition˜3.1
@@ A.1 Assessment of the scale to use
Suppose $M_{2}$ is another mesh that is a scaled version of $M_{1}$ by a scaling factor $a$. Then, according to (Bronstein & Kokkinos, 2010):
@@
The heat kernel can be computed as per (Sun et al., 2009):
@@ Appendix C Implementation Details
C.1 Protein representation details
Surface representation
To generate the surface representation $\mathcal{S}_{\mathbf{P}}$ of the protein $\mathbf{P}$, our initial step involves computing the protein surface using MSMS (Sanner et al., 1996). The resulting meshes are usually quite large, but in the rare cases of small proteins, we incrementally increase sampling density until we achieve a minimum number of 256 vertices. Then, we employ quadratic decimation (Garland & Heckbert, 1997), to embed our surfaces into coarser and more compact meshes, again ensuring a minimum vertex count. Coarsening the mesh has an impact on computational efficiency and on performance, assessed in Sections 3.5 and 4. Finally, we ensure mesh quality by removing non-manifold edges, duplicated or degenerate vertices and faces, as well as small disconnected components. We compute several surface features: the Gaussian and mean curvature, as well as the shape index at each vertex. We also include the normal as well as the heat kernel signature (Sun et al., 2009) at each point.
Graph representation
In addition to the surface $\mathcal{S}_{\mathbf{P}}$, we construct a graph representation $\mathcal{G}_{\mathbf{P}}=(\mathcal{V}_{g},\mathcal{E}_{g})$. To ensure consistency and fairness in comparison within the Atom3d benchmark, we follow their conventions: each node in the graph corresponds to an atom within the protein and edges $\mathcal{E}_{g}$ are defined between pairs of atoms that are within a 4.5 angstrom radius cutoff. The only initial node features are one-hot encoding of atom types.
Bipartite graph construction
[…]
C.2 Model architecture details
@@
However, the sequential approach does not simultaneously leverage the distinct scales (local and global) offered by graph and surface processing. To overcome this, the bipartite approach processes graph and surface features concurrently in separate encoders, and merge them with a message passing incorporating learnable parameters, based on the equation: $\mathcal{X}_{out}=\alpha\mathcal{X}+\text{MP}_{gs}(\mathcal{H})+\text{MP}_{sg}(\mathcal{H})$. Those architectures are illustrated in Figure˜4.
Supplementary Figure 4: Illustration of our approach integrating surface and graph information.
@@ D.1 Complexity analysis
[…]
@@ D.2 Memory and batching
In addition to individual objects being large, the large variance in the object size (see Figure˜8) also prevents us from using large batch sizes. Indeed, due to the random composition of the batch, some unlucky batches contain several particularly large proteins, causing spikes in memory usage and out-of-memory errors. We validate in Figure˜8 that using coarsened surfaces, we stay in the linear dependency on the number of vertices for the hidden sizes at stake.
@@
[…]
This is illustrated in Figure˜9, for which we tracked the memory consumption of our model through one epoch of learning. Observe the memory peaks that can result in over 20% excess memory consumption, keeping in mind that this is only one run over one epoch. When training several models for hours, the probability that one outlier results in a 50% increase forces user caution.
Supplementary Figure 9: Memory consumption monitoring through an epoch, for different batch sizes.